OpenStudy (moonlitfate):

Use trigonometric identities to transform the left side of the equation into the right side (0 < θ < π/2). sin^2θ − cos^2θ = sin^2θ - (______) = 2sin^2θ - 1

hartnn (hartnn):

cos^2θ can be written as 1-sin^2 θ

hartnn (hartnn):

because cos^2 θ+sin^2 θ = 1

OpenStudy (moonlitfate):

So, \[\sin ^{2}\theta - (1-\sin^{2} \theta) = 2\sin^{2} \theta -1 \]

hartnn (hartnn):

yes,thats correct :)